Use the method of substitution to solve the system of linear equations.
3a=4-b
3a=3-b

Answer by MathLover1(20849)


3a=4-b.......eq.1
3a=3-b.......eq.2
........................substitute 4-b for 3a+ from eq.1 into eq.2
4-b=3-b.......eq.2
b-b=3-4
0=3-4 which is not true, the two lines are parallel and will never intersect; so, this system is inconsistent